# Annabelle Michael Carrell    Albert Gong    Abhishek Shetty    Raaz Dwivedi    Lester Mackey

### v2

@@ 4.2 Thinning attention in practice
To gauge the practical effectiveness of Alg. 1, we recreate the benchmark Tokens-To-Token Vision Transformer (T2T-ViT) experiment of Zandieh et al. (2023). In this experiment, attention approximations are scored on their ImageNet classification accuracy and computational expense when used as drop-in replacements for the two most expensive attention layers in a pretrained T2T-ViT neural network (Yuan et al., 2021). Using the exact implementations and settings provided by Zandieh et al. (2023), we benchmark our PyTorch implementation of Thinformer against exact attention and four leading attention approximations: Performer (Choromanski et al., 2021), Reformer (Kitaev et al., 2020), ScatterBrain (Chen et al., 2021), and KDEformer. In Tab. 3, we find that Thinformer provides the highest Top-$1$ accuracy on the ImageNet 2012 validation set (Russakovsky et al., 2015), while running in time comparable to the fastest alternative. The final attention call of Thinformer can also be combined with optimized attention implementations like FlashAttention (Dao et al., 2022; Dao, 2024) to further reduce the time and memory footprint. We provide PyTorch code replicating this experiment at https://github.com/microsoft/thinformer and supplementary experiment details in Sec. L.1.
@@ 6.1 Low-rank analysis of Compress Then Test
$\textstyle\widehat{\mathbsf{R}}_{\mathbf{k}}^{2}\triangleq\log(\frac{n}{s})\log(\frac{n}{\widetilde{\beta}})\,\cdot$ (41)
@@
Remark 1.
Combining the argument of App. J with the distinct analysis of Dwivedi & Mackey (2022, Thm. 2, Prop. 3) would yield the worse dimension dependence
$\mathbsf{R}_{\mathbf{k}_{\textup{deep}}}=O(\log^{\frac{3d^{\prime}}{4}+2}(\frac{n}{\widetilde{\beta}})).$ (46)
@@ B.1.2 Sub-Gaussianity of subsampling
$|\boldsymbol{u}^{\top}\mathbf{K}\boldsymbol{e}_{J_{i}}|\leq\sqrt{\boldsymbol{u}^{\top}\mathbf{K}\boldsymbol{u}}\sqrt{\boldsymbol{e}_{J_{i}}^{\top}\mathbf{K}\boldsymbol{e}_{J_{i}}}\leq\sqrt{\|{\mathbf{K}}\|_{\max}}\sqrt{\boldsymbol{u}^{\top}\mathbf{K}\boldsymbol{u}}\quad\text{surely}\quad$ (65)
@@ C.1 Proof of MMD bound 12
$\|{\mathbf{V}_{-r}^{\top}\boldsymbol{\Phi}^{\top}\boldsymbol{w}}\|_{2}^{2}\leq\lambda_{r+1}(\frac{1}{n_{\textup{out}}}-\frac{1}{n})\quad\text{surely,}\quad$, (111)
@@
$\|{\boldsymbol{w}}\|_{2}^{2}=\boldsymbol{p}_{\textup{in}}^{\top}\boldsymbol{p}_{\textup{in}}+\boldsymbol{p}_{\textup{out}}^{\top}\boldsymbol{p}_{\textup{out}}-2\boldsymbol{p}_{\textup{in}}^{\top}\boldsymbol{p}_{\textup{out}}=\frac{n}{n^{2}}+\frac{n_{\textup{out}}}{n_{\textup{out}}^{2}}-\frac{2n_{\textup{out}}}{nn_{\textup{out}}}=\frac{1}{n_{\textup{out}}}-\frac{1}{n},$ (133)
@@ C.2 Proof of kernel max seminorm bound 13
The triangle inequality and the covering property 136 together imply that, surely,
@@ C.3 Proof of Lipschitz kernel max seminorm bound 15
Since, by our Lipschitz assumption,
$\|{\mathbf{k}(\boldsymbol{z},\cdot)-\mathbf{k}(\boldsymbol{z}^{\prime},\cdot)}\|_{\mathbf{k}}^{2}=\mathbf{k}(\boldsymbol{z},\boldsymbol{z})-\mathbf{k}(\boldsymbol{z},\boldsymbol{z}^{\prime})+\mathbf{k}(\boldsymbol{z}^{\prime},\boldsymbol{z}^{\prime})-\mathbf{k}(\boldsymbol{z}^{\prime},\boldsymbol{z})\leq\min(4\max_{\boldsymbol{z}\in\mathcal{Z}}\mathbf{k}(\boldsymbol{z},\boldsymbol{z}),2L_{\mathbf{k}}\|{\boldsymbol{z}-\boldsymbol{z}^{\prime}}\|_{2}),$ (169)
@@
$\tilde{A}=\sqrt{\log(\frac{2n(\log(n/2)+1)}{\delta})\mathopen{}\mathclose{{}\left[e^{2}r+e\log(\frac{n}{\delta^{\prime}})}\right]+4}.$ (213)
@@
Define $\mathbf{X}=[\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}]^{\top}$ and $\mathcal{X}=\{\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}\}$, let $c=2\max_{i\in[n]}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}$ for $\bar{\boldsymbol{x}}^{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}$, and set $r=\mathrm{rank}_{c/\sqrt{n}}(\mathbf{X})$. For any $j\in[n]$, we can write
$\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}=\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}-\sum_{i=1}^{j}\mathbf{1}\mathopen{}\mathclose{{}\left[{\boldsymbol{x}_{i}\in\mathcal{X}_{\textup{out}}^{j}}}\right]\boldsymbol{x}_{i}^{k}}\|_{2}=2j\|{\mathbf{X}^{\top}(\boldsymbol{p}_{\textup{in}}^{j}-\boldsymbol{p}_{\textup{out}}^{j})}\|_{2}=2j\operatorname{MMD}_{\mathbf{X}\mathbf{X}^{\top}}(\boldsymbol{p}_{\textup{in}}^{j},\boldsymbol{p}_{\textup{out}}^{j})$ (214)
where $\boldsymbol{p}_{\textup{in}}^{j}$ and $\boldsymbol{p}_{\textup{out}}^{j}$ are the empirical distributions over $\mathcal{X}_{\textup{in}}^{j}=(\boldsymbol{x}_{i}^{k})_{i=1}^{j}$ and $\mathcal{X}_{\textup{out}}^{j}=\{\boldsymbol{x}_{i}\in\mathcal{X}_{\textup{out}}:i\in[j]\}$.
Since LKH$(\delta)$ is an online algorithm that assigns signs $(\epsilon_{i}^{k},\epsilon_{i+1}^{k}=1-\epsilon_{i}^{k})$ to the points $(\boldsymbol{x}_{i}^{k},\boldsymbol{x}_{i+1}^{k})$ sequentially, we can view $\mathcal{X}_{\textup{out}}^{j}$ as the output of LKH$(\delta)$ applied to $\mathcal{X}_{\textup{in}}^{j}$ with $n_{\textup{out}}=\frac{j}{2}$ and the linear kernel $\mathbf{k}(\boldsymbol{x},\boldsymbol{y})=\langle{\boldsymbol{x}},{\boldsymbol{y}}\rangle$ for each $j\in[n]$. Therefore, we may invoke the established LKH$(\delta)$ sub-Gaussian constants $\nu_{j}$ of Prop. B.3, Thm. 1, the union bound, and the definition of $\epsilon$-rank (Def. 4) to deduce that
$\max_{j\in[n]}\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}^{2}$ $\leq\max_{j\in[n]}4j^{2}\nu_{j}^{2}\mathopen{}\mathclose{{}\left[e^{2}r+e\log(\frac{n}{\delta^{\prime}})}\right]+\sigma_{r+1}(\mathbf{X})^{2}\frac{4j^{2}}{j}\leq c^{2}\tilde{A}^{2}$ (215)
@@ Appendix H KT-Compress$(\delta)$
Thm. 4 follows from the following more detailed statement, proved in Sec. I.1, by taking
$\widehat{\mathbsf{R}}_{\mathbf{k}}^{2}\triangleq\mathbsf{R}^{2}(\mathbf{K},n_{\textup{in}},\frac{\widetilde{\beta}}{20s_{n}},\mathfrak{g})+\mathbsf{R}^{2}(\mathbf{K}^{\prime},n_{\textup{in}},\frac{\widetilde{\beta}}{20s_{n}},\mathfrak{g}).$ (216)
@@
The experiment of Sec. 4.2 was carried out using Python 3.12.5, PyTorch 2.6.0 (Paszke et al., 2019), and an Ubuntu 20.04 server with 8 CPU cores (Intel(R) Xeon(R) Silver 4316 CPU @ 2.30GHz), 100 GB RAM, and a single Nvidia A6000 GPU (48 GB memory, CUDA 12.1, driver version 530.30.02). For reference, attention layer 1 has $(n,d)=(3136,64)$ and attention layer 2 has $(n,d)=(784,64)$. For each layer and each of the first $50$ ImageNet 2012 validation set batches of size $64$, we measured the time required to complete a forward pass through the layer using CUDA events following $10$ warm-up batches to initialize the GPU. Tab. L.1 provides the hyperparameter settings for each attention approximation in Tab. 3. The settings and implementations for all methods other than Thinformer were provided by Zandieh et al. (2023), and our experiment code builds on their open-source repository https://github.com/majid-daliri/kdeformer.
@@
The settings and implementation for both random reshuffling (RR) and CD-GraB: Greedy were those used in the original logistic regression on mortgage application experiment of Cooper et al. (2023). Our experiment code builds on the open-source CD-GraB repository https://github.com/GarlGuo/CD-GraB.
@@
The experiment of Sec. 6.2 was carried out using Python 3.10.15, PyTorch 2.5.0, and a Rocky Linux 8.9 server with an AMD EPYC 9454 48-Core Processor, 100 GB RAM, and a single Nvidia H100 GPU (80 GB memory, CUDA 12.5, driver version 555.42.02). Each test is run with replication count $\mathcal{B}=100$, nominal level $\alpha=0.05$, and failure probability $\delta=0.5$. Runtime measurements exclude the time required to train the neural network $\phi$. Our experiment code builds on the open-source deep kernel testing (https://github.com/fengliu90/DK-for-TST) and Compress Then Test (https://github.com/microsoft/goodpoints) repositories.

### v4

@@ 4.2 Thinning attention in practice
To gauge the practical effectiveness of Alg. 1, we recreate the benchmark Tokens-To-Token Vision Transformer (T2T-ViT) experiment of Zandieh et al. (2023). In this experiment, attention approximations are scored on their ImageNet classification accuracy and computational expense when used as drop-in replacements for the two most expensive attention layers in a pretrained T2T-ViT neural network (Yuan et al., 2021). Using the exact implementations and settings provided by Zandieh et al. (2023), we benchmark our PyTorch implementation of Thinformer against exact attention and four leading attention approximations: Performer (Choromanski et al., 2021), Reformer (Kitaev et al., 2020), ScatterBrain (Chen et al., 2021), and KDEformer. In Tab. 3, we find that Thinformer provides the highest Top-$1$ accuracy on the ImageNet 2012 validation set (Russakovsky et al., 2015), while running faster than all of the alternatives. The final attention call of Thinformer can also be combined with optimized attention implementations like FlashAttention (Dao et al., 2022; Dao, 2024) to further reduce the time and memory footprint. We provide PyTorch code replicating this experiment at https://github.com/microsoft/thinformer and supplementary experiment details in Sec. L.1.
@@ 6.1 Low-rank analysis of Compress Then Test
$\textstyle\widehat{\mathbsf{R}}_{\mathbf{k}}^{2}\triangleq\log(\frac{m+n}{s})\log(\frac{n}{\widetilde{\beta}})\,\cdot$ (41)
@@
Cors. 3 and 4 follow from explicitly bounding the eigenvalues of the generated deep kernel matrices as in 17 and 20. One could alternatively bound the compression error of KT-Compress$(\delta)$ using the covering number approach of Dwivedi & Mackey (2022, Thm. 2, Prop. 3). In the setting of Cor. 3, the argument of App. J combined with this distinct analysis would yield an alternative error inflation factor $\mathbsf{\tilde{R}}_{\mathbf{k}_{\textup{deep}}}/2^{\mathfrak{g}}$ with worse dimension dependence,
$\mathbsf{\tilde{R}}_{\mathbf{k}_{\textup{deep}}}=\Theta(\log^{\frac{3d^{\prime}}{4}+2}(\frac{n}{\widetilde{\beta}})),$ (47)
and without known adaptivity to an intrinsic manifold dimension.
@@ 6.2 Powerful deep kernel testing in near-linear time
Acknowledgments
We thank Insu Han, A. Feder Cooper, and Wentao Guo for their assistance with their code bases and datasets.
@@ B.1.2 Sub-Gaussianity of subsampling
$|\boldsymbol{u}^{\top}\mathbf{K}\boldsymbol{e}_{J_{i}}|\leq\sqrt{\boldsymbol{u}^{\top}\mathbf{K}\boldsymbol{u}}\sqrt{\boldsymbol{e}_{J_{i}}^{\top}\mathbf{K}\boldsymbol{e}_{J_{i}}}\leq\sqrt{\|{\mathbf{K}}\|_{\max}}\sqrt{\boldsymbol{u}^{\top}\mathbf{K}\boldsymbol{u}}\quad\text{with probability $1$}\quad$ (65)
@@ C.1 Proof of MMD bound 12
$\mathbb{P}(\|{\mathbf{V}_{-r}^{\top}\boldsymbol{\Phi}^{\top}\boldsymbol{w}}\|_{2}^{2}\leq\lambda_{r+1}(\frac{1}{n_{\textup{out}}}-\frac{1}{n}))=1,$ (111)
@@
$\|{\boldsymbol{w}}\|_{2}^{2}=\boldsymbol{p}_{\textup{in}}^{\top}\boldsymbol{p}_{\textup{in}}+\boldsymbol{p}_{\textup{out}}^{\top}\boldsymbol{p}_{\textup{out}}-2\boldsymbol{p}_{\textup{in}}^{\top}\boldsymbol{p}_{\textup{out}}=\frac{n_{\textup{in}}}{n_{\textup{in}}^{2}}+\frac{n_{\textup{out}}}{n_{\textup{out}}^{2}}-\frac{2n_{\textup{out}}}{n_{\textup{in}}n_{\textup{out}}}=\frac{1}{n_{\textup{out}}}-\frac{1}{n_{\textup{in}}},$ (133)
@@ C.2 Proof of kernel max seminorm bound 13
The triangle inequality and the covering property 136 together imply that, with probability $1$,
@@ C.3 Proof of Lipschitz kernel max seminorm bound 15
Moreover, by our Lipschitz assumption,
$\|{\mathbf{k}(\boldsymbol{z},\cdot)-\mathbf{k}(\boldsymbol{z}^{\prime},\cdot)}\|_{\mathbf{k}}^{2}=\mathbf{k}(\boldsymbol{z},\boldsymbol{z})-\mathbf{k}(\boldsymbol{z},\boldsymbol{z}^{\prime})+\mathbf{k}(\boldsymbol{z}^{\prime},\boldsymbol{z}^{\prime})-\mathbf{k}(\boldsymbol{z}^{\prime},\boldsymbol{z})\leq\min(4\max_{\boldsymbol{z}\in\mathcal{Z}}\mathbf{k}(\boldsymbol{z},\boldsymbol{z}),2L_{\mathbf{k}}\|{\boldsymbol{z}-\boldsymbol{z}^{\prime}}\|_{2}).$ (169)
@@
$\tilde{A}=\sqrt{\log(\frac{2n(\log(n/2)+1)}{\delta})\mathopen{}\mathclose{{}\left[e^{2}\,\mathrm{rank}_{\epsilon_{k}}(\mathbf{X}^{k})+e\log(\frac{n}{\delta^{\prime}})}\right]+1}$ (213)
for $\mathbf{X}^{k}\triangleq[\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}]^{\top}$, $\epsilon_{k}\triangleq{\max_{i\in[n]}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}}/{\sqrt{n}},$ and $\bar{\boldsymbol{x}}^{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}.$
@@
Define $\mathcal{X}^{k}=\{\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}\}$, $c=2\max_{i\in[n]}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}$, and $r=\mathrm{rank}_{\epsilon_{k}}(\mathbf{X}^{k})$. For any $j\in[n]$, we can write
$\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}=\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}-\sum_{i=1}^{j}\mathbf{1}\mathopen{}\mathclose{{}\left[{\boldsymbol{x}_{i}^{k}\in\mathcal{X}_{\textup{out},j}^{k}}}\right]\boldsymbol{x}_{i}^{k}}\|_{2}=2j\|{(\mathbf{X}^{k})^{\top}(\boldsymbol{p}_{\textup{in}}^{j}-\boldsymbol{p}_{\textup{out}}^{j})}\|_{2}=2j\operatorname{MMD}_{\mathbf{X}^{k}(\mathbf{X}^{k})^{\top}}(\boldsymbol{p}_{\textup{in}}^{j},\boldsymbol{p}_{\textup{out}}^{j})$ (214)
where $\boldsymbol{p}_{\textup{in}}^{j}$ and $\boldsymbol{p}_{\textup{out}}^{j}$ are the empirical distributions over $\mathcal{X}_{\textup{in},j}^{k}=(\boldsymbol{x}_{i}^{k})_{i=1}^{j}$ and $\mathcal{X}_{\textup{out},j}^{k}=\{\boldsymbol{x}_{i}^{k}\in\mathcal{X}_{\textup{out}}^{k}:i\in[j]\}$.
Since LKH$(\delta)$ is an online algorithm that assigns signs $(\epsilon_{i}^{k},\epsilon_{i+1}^{k}=1-\epsilon_{i}^{k})$ to the points $(\boldsymbol{x}_{i}^{k},\boldsymbol{x}_{i+1}^{k})$ sequentially, we can view $\mathcal{X}_{\textup{out},j}^{k}$ as the output of LKH$(\delta)$ applied to $\mathcal{X}_{\textup{in},j}^{k}$ with $n_{\textup{out}}=\frac{j}{2}$ and the linear kernel $\mathbf{k}(\boldsymbol{x},\boldsymbol{y})=\langle{\boldsymbol{x}},{\boldsymbol{y}}\rangle$ for each $j\in[n]$. Therefore, we may invoke the established LKH$(\delta)$ sub-Gaussian constants $\nu_{j}$ of Prop. B.3, Thm. 1, the union bound, and the definition of $\epsilon$-rank (Def. 4) to deduce that
$\max_{j\in[n]}\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}^{2}$ $\leq\max_{j\in[n]}4j^{2}\nu_{j}^{2}\mathopen{}\mathclose{{}\left[e^{2}r+e\log(\frac{n}{\delta^{\prime}})}\right]+\sigma_{r+1}(\mathbf{X}^{k})^{2}\frac{4j^{2}}{j}\leq c^{2}\tilde{A}^{2}$ (215)
@@ Appendix H KT-Compress$(\delta)$
Thm. 4 follows from the following more detailed statement, proved in Sec. I.1 as
$\mathbsf{R}_{\mathbf{K}}^{2}(n_{\textup{in}},\frac{\widetilde{\beta}}{20s_{n}},\mathfrak{g})+\mathbsf{R}_{\mathbf{K}^{\prime}}^{2}(n_{\textup{in}},\frac{\widetilde{\beta}}{20s_{n}},\mathfrak{g})=O(\widehat{\mathbsf{R}}_{\mathbf{k}}^{2}).$ (216)
@@
The experiment of Sec. 4.2 was carried out using Python 3.12.9, PyTorch 2.8.0.dev20250407+cu128 (Paszke et al., 2019), and an Ubuntu 22.04.5 LTS server with an AMD EPYC 7V13 64-Core Processor, 220 GB RAM, and a single NVIDIA A100 GPU (80 GB memory, CUDA 12.8, driver version 570.124.04). For reference, attention layer 1 has $(n,d)=(3136,64)$ and attention layer 2 has $(n,d)=(784,64)$. For each layer and each of the first $50$ ImageNet 2012 validation set batches of size $64$, we measured the time required to complete a forward pass through the layer using CUDA events following $10$ warm-up batches to initialize the GPU. Tab. L.1 provides the hyperparameter settings for each attention approximation in Tab. 3. The settings and implementations for all methods other than Thinformer were provided by Zandieh et al. (2023), and our experiment code builds on their open-source repository https://github.com/majid-daliri/kdeformer.
@@
The settings and implementation for both random reshuffling (RR) and CD-GraB: Greedy were those used in the original logistic regression on mortgage application experiment of Cooper et al. (2023). Our experiment code builds on the open-source CD-GraB repository https://github.com/GarlGuo/CD-GraB. As in Cooper et al. (2023), optimization was carried out with a learning rate of $\alpha=0.01$, datapoints were loaded in batches of size $16$, and stochastic gradients were reordered for each datapoint individually.
@@
The experiment of Sec. 6.2 was carried out using Python 3.10.15, PyTorch 2.5.0, and a Rocky Linux 8.9 server with an AMD EPYC 9454 48-Core Processor, 100 GB RAM, and a single NVIDIA H100 GPU (80 GB memory, CUDA 12.5, driver version 555.42.02). Each test is run with replication count $\mathcal{B}=100$, nominal level $\alpha=0.05$, and failure probability $\delta=0.5$. The neural network $\phi$ was trained exactly as in Liu et al. (2020) (with learning rate $5\times 10^{-5}$ and batch size equal to the full training sample size), and runtime measurements exclude the time required to train $\phi$. Our experiment code builds on the open-source deep kernel testing (https://github.com/fengliu90/DK-for-TST) and Compress Then Test (https://github.com/microsoft/goodpoints) repositories.